I'm having some problem with the following:
Mary starts at a point 26 miles from school & runs toward school at 6 miles per hour. Mike starts from school 1.3 hours later & runs toward Mary at 8 miles per hour.
A) Let x be the number of hours Mary has been running. Write expressions for the distance she has run & for her distance from school.
B) In terms of x, how many hours has Mike been running? How far is he from school?
C) How long has Mary been running when she & Mike meet? (write equation & solve)
D) How far from school do they meet?

Answer by scott8148(6628)

(A) d = r * t = 6x
___ distance from school = 26 - 6x

(B) t = x - 1.3
___ distance from school = 8(x - 1.3))

(C) 26 - 6x = 8(x - 1.3))
___ 26 - 6x = 8x - 10.4
___ 36.4 = 14x
___ 36.4 / 14 = x

(D) 26 - (6(36.4 / 14))